OpenStudy (peace420):

The first term of an arithmetic sequence is 4 and the tenth term is 67. What is the common difference? Show your work or explain your answer.

OpenStudy (tkhunny):

It's an ARITHMETIC sequence. This is the big item in the question. 1) a 2) a+d 3) a + 2d 4) a + 3d ... 10) You tell me.

satellite73 (satellite73):

you got this or no?

OpenStudy (peace420):

no

satellite73 (satellite73):

tenth one is \[4+9d\] which you are told is 67, set \[4+9d=37\] and solve for \(d\) in two steps

OpenStudy (sshayer):

tn=a+(n-1)d

satellite73 (satellite73):

oops i meant \[4+9d=67\] doh

satellite73 (satellite73):

subtract 4, divide by 9 to get your answer

OpenStudy (vuriffy):

Did you get your answer? 4 + 9d = 67 (Subtract by 4) 9d = 63 (divide by 9) d = ?

OpenStudy (peace420):

Thank you everyone

OpenStudy (peace420):

!!